eljaja:

Can someone simplify this for me real quick? (sinθ-cosθ) - (sinθ + cosθ)^2 it doesn't seem to work in any of my calculators

justasoftie:

I can't help you with this, but @Vocaloid may be able to.

nuts:

well, why don't you try doing it by hand? expand the second term

bananas:

\[Sin - Cos - Sin^2 - 2SinCos - Cos^2 =\]\[Sin - Cos - 2SinCos - Sin^2 - Cos^2 = Sin - Cos - 2SinCos-(Sin^2+Cos^2)=\]\[Sin-Cos-2SinCos-1\]did I miss anything nuts?

nuts:

i think she typed the question wrong and didn't put in the ^2 in the first expression but i'm not 100% sure

nuts:

2 sin cos can be written as sin (2 th)

eljaja:

no.... I didn't type anything wrong. Thats how the question looks... im not sure if ^2 is for both expressions but it is for the second one.

nuts:

aight then bananas is correct prolly

bananas:

Then going off what @nuts said,\[Sin−Cos−2SinCos−1=Sin\theta-Cos\theta-Sin2\theta-1\]

bananas:

Throw in all the thetas that I didn't bother typing obviously